OpenStudy (ria23):

Solve for x e^xe^(x+1)=1

OpenStudy (ria23):

\[e^{x}e^{(x+1)}=1\]

hartnn (hartnn):

you can start by dividing e^x on both sides

hartnn (hartnn):

then compare the exponents on both sides :)

OpenStudy (australopithecus):

wouldn't it be better to just write it as e^(2x+1) = 1 then take the ln of both sides, the exploit the rule ln(a^b) = bln(a)

hartnn (hartnn):

\(\Large \dfrac{1}{a^b} = a^{-b}\)

OpenStudy (ria23):

Would that cancel e^x on the left ad put it right next to the 1? \[e^{(x+1)}=1e^{x}\]

hartnn (hartnn):

thats a good approach too Austral :)

hartnn (hartnn):

\(\large e^{x+1}= e^{-x}\)

hartnn (hartnn):

x+1 =-x

OpenStudy (solomonzelman):

\(\large\color{black}{ e^{x}e^{x+1}=1 }\) \(\large\color{black}{ e^{x+x+1}=1 }\) \(\large\color{black}{ e^{2x+1}=1 }\) \(\large\color{black}{ e^{2x+1}=e^{\ln(1)} }\) \(\large\color{black}{ 2x+1=\ln(1) }\) and we all know what \(\large\color{black}{ \ln(1) }\) is.

hartnn (hartnn):

because \(\Large \dfrac{1}{e^x} = e^{-x}\)

hartnn (hartnn):

my method got away with the log part .-.

OpenStudy (solomonzelman):

there is actually no "lol" part, or not that I would say that my solution involves logarithms.

OpenStudy (solomonzelman):

Opps, log part, not lol part.

hartnn (hartnn):

lol

hartnn (hartnn):

ria, confused? got it ? start over?

OpenStudy (ria23):

Uhm... ln(1)=0 So... Wait I have a couple questions... :D Yhu got 2x+1 because yhu combined the like terms right?

hartnn (hartnn):

\( \Large a^ma^n = a^{m+n}\) so yeah x+ (x+1) combining like terms gives 2x+1

OpenStudy (solomonzelman):

yes, added exponents of e, to get 2x+1

OpenStudy (ria23):

And ln=0, from there, would 2x+1 be my answer, or would I need to solve it from there like I did with the last one. x=-2?

hartnn (hartnn):

yhu need to solve 2x+1= 0 and x is NOT -2

OpenStudy (ria23):

.-. -1/2?

hartnn (hartnn):

yheshhhh

hartnn (hartnn):

i mean thats correct :) :P

OpenStudy (ria23):

^o^ Yea!

OpenStudy (ria23):

Thank yhu both ^u^ I'm gonna medal Solomon cuz he didn't get a medal. Thank yhu again so much Hartnn ^.^ Yhur really helpful. n~n

hartnn (hartnn):

There should be one for everyone ;) most welcome ^_^

OpenStudy (solomonzelman):

I did pretty much the same as taking the natural log of both sides, but I used a rule \(\Large\color{black}{ \color{red}{a}=e^{\ln(\color{red}{a})} }\) you can prove this rule, if you lo=ike by taking the natural log of both sides. \(\Large\color{black}{ \ln\color{red}{a}=\ln \left(\begin{matrix} e^{\ln(\color{red}{a})} \\ \end{matrix}\right) }\) taking the exponent out, \(\Large\color{black}{\ln( \color{red}{a})={\ln(\color{red}{a})}\ln(e) }\) we know that, ( \(\Large\color{black}{\ln(e)=\log_ee=1 }\) ) \(\Large\color{black}{\ln( \color{red}{a})={\ln(\color{red}{a})} }\) \(\Large\color{black}{ \color{red}{a}={\color{red}{a}} }\)

OpenStudy (solomonzelman):

this rule helped me a lot in all of my courses:)

OpenStudy (ria23):

:O Ahh... Dude... Thank yhu. n_n

OpenStudy (solomonzelman):

np